OpenStudy (anonymous):

pythagorean trile 1 2 400

OpenStudy (anonymous):

4-12-20 is not a Pythagorean trile because 4^2 + 12^2 is NOT equal to 20^2 16 + 144 is NOT equal to 400 160 is NOT equal to 400

OpenStudy (anonymous):

All other triplets are Pytagorean triplets.

OpenStudy (anonymous):

The second question: 8^2 + 12^2 = x^2 64 + 144 = x^2 208 = x^2 x = sqrt(208)

OpenStudy (anonymous):

Thank you.